A household consuming a bundle of goods X and Y such that MUX = 15 and MUY = 3 when the prices are such that pX = 5 and pY = 1 is maximizing its “utility”.
1
Expert's answer
2014-10-28T14:29:53-0400
MUX = 15, MUY = 3, Px = 5, Py = 1
The utility is maximized, when MUx/Px = MUy/Py
In our case MUx/Px = 15/5 = 3 and MUy/Py = 3/1 = 3, so the utility is
maximized.